three partners among a firm divide a profit of rs.67,200. among them selves in the ratio 2:3:7 find the share of each one of them?


1
Expert's answer
2021-09-28T01:12:43-0400


Total = "2+3+7=12"


So the partners received the following amount


"\frac{2}{12}×67200=10120"


"\frac{3}{12}×67200=15180"


"\frac{7}{12}×67200=35420"